OpenStudy (anonymous):

Given f(x) = 6x2 – x – 12 and g(x) = 2x – 3, find the function (fg)(x).

OpenStudy (anonymous):

@shivamgautam can you help with this as well?

OpenStudy (anonymous):

u mean we have to find f(x) right?

OpenStudy (anonymous):

@WaVVessux

OpenStudy (anonymous):

(fg)(x)

OpenStudy (anonymous):

if that is right then first of all find x from g(x)=2x-3 as x=3/2 now put value of x in f(x): 6(3/2)^2 - 3/2 - 12

OpenStudy (anonymous):

okay is that the answer?

OpenStudy (anonymous):

k its easy (fg)(x) means f(x)*g(x)

OpenStudy (anonymous):

alright, so now I solve then?

OpenStudy (anonymous):

now continue with above process f(x) = 6.9/4 - 3/2 -12 = 54/4 - 3/2 -12 = 27/2 - 3/2 -12 = 24/ 2 - 12 = 0

OpenStudy (anonymous):

Oh okay I got 0 but I didn't think it was right

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

...I'm lost :/

OpenStudy (anonymous):

wait a min i will get back to shortly

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

k here we have to find (fg)(x) it means (fx)*(gx)

OpenStudy (anonymous):

so first multiple both functions

OpenStudy (anonymous):

we will get 12 x^3-20 x^2-21 x+36

OpenStudy (anonymous):

now we can solve this using cube formula

OpenStudy (anonymous):

should i give u all steps!

OpenStudy (jhannybean):

Allyou have to do is multiply these two functions.. :) Omg the cube formula.. -_-

OpenStudy (anonymous):

and answer will be 3/2 and -4/3

OpenStudy (anonymous):

but, I need it in polynomial standard form...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

we can use alternate forms to solve this 12 x^3-20 x^2-21 x+36 alternative form is (3-2x)^2(3x+4)

OpenStudy (anonymous):

now we can solve them separately as 3x+4=0 9+4x^2-12x

OpenStudy (anonymous):

And that's it? :D

OpenStudy (anonymous):

(1) 3x+4=0 3x=-4 x=-4/3 (2) 4x^2-12x+9=0 so after solving it from quadratic formula: x=3/2

OpenStudy (anonymous):

and that it!!! :)

OpenStudy (anonymous):

For the standard polynomial form?

OpenStudy (anonymous):

and we could have solved it from the first answer i told u by find x from 3/2 after putting value of x in f(x) we get zero that means 3/2 satisfies the f(x) therefore 3/2 is the answer!! :)

OpenStudy (anonymous):

@McLove

OpenStudy (anonymous):

12 x^3-20 x^2-21 x+36 (this will be the answer for standard polynomial form)

OpenStudy (anonymous):

Much gratitude.

OpenStudy (jhannybean):

Hm..

OpenStudy (anonymous):

??

OpenStudy (anonymous):

@Jhannybean first take look at the question we were given it is f(x)=6x2 – x – 12 and g(x) = 2x-3

OpenStudy (jhannybean):

f x g = (6x^2 - x - 12)(2x - 3) = 12x^3 - 2x^2 -24x -18x^2 +3x+36 12x^3 -20x^2 -21x +36

OpenStudy (anonymous):

first alternative can be found by above info so just simply f(x) and ur alternative form will be there

OpenStudy (anonymous):

did u get it @Jhannybean

OpenStudy (jhannybean):

Hm... first alternative? let me scroll up and see...

OpenStudy (anonymous):

oh sorry i mean alternative

OpenStudy (jhannybean):

Yeah... that's where im getting stuck. "alternative form is (3-2x)^2(3x+4)" how did you derive this equation to standard form?

OpenStudy (jhannybean):

I mean....break it down into those....

OpenStudy (anonymous):

we first break f(x) and will get (2x-3)(4+3x)

OpenStudy (anonymous):

now as we have g(x) to answer will be 12 x^3-20 x^2-21 x+36 = (2x-3)(4+3x)(2x-3)

OpenStudy (jhannybean):

ohhhh!!! HAHA got it.

OpenStudy (anonymous):

as we have two 2x-3 it will be (2x-3)^2(4+3x)

OpenStudy (jhannybean):

SMART. i don't know why i didnt see that.... lol.

OpenStudy (anonymous):

hehe :) u forgot g(x)

OpenStudy (jhannybean):

yep.

OpenStudy (anonymous):

bye

OpenStudy (jhannybean):

i was trying to work back from the expanded form and i was like.... what!?